OpenStudy (anonymous):

If y varies inversely as x, find the constant of variation if y = 36 when x = 9. A. 27 B. 14 C. 324 D. 4

OpenStudy (paki):

y = k/x... 36 = k/9 ==> k = 324 ... @musa19965